# TimeMachine: A Time Series is Worth 4 Mambas for Long-term Forecasting

### v1

@@ 1 Introduction
Capturing long-term dependencies, which are generally abundant in MTS data, is pivotal to LTSF. While linear models such as DLineaer [33] and TiDE [6] achieve competitive performance with linear complexity and scalability, with accuracy on par with Transformer-based models, they usually rely on MLPs and linear projections that may not well capture long-range correlations [4]. Transformer-based models such as iTransformer [35], PatchTST [23], and Crossformer [35] have a strong ability to capture long-range dependencies and superior performance in LTSF accuracy, thanks to the self-attention mechanisms in Transformers [28]. However, they typically suffer from the quadratic complexity [6], limiting their scalability and applicability, e.g., in edge computing.
@@
The under-utilization of SSMs in LTSF can be attributed to two main reasons. First, highly content- and context-selective SSMs have only been recently developed [11]. Second, and more importantly, effectively representing the context in time series data remains a challenge. Many Transformer-based models, such as Autoformer [29] and Informer [36], regard each time point as a token in a sequence, while more recent models like PatchTST [24] and iTransformer [21] leverage patches of the time series as tokens. However, our empirical experiments on real-world MTS data suggest that directly utilizing SSMs for LTSF by using either time points or patches as tokens could hardly achieve performance comparable to Transformer-based models. Considering the particular characteristics of MTS data, it is essential to extract more salient contextual cues tailored to SSMs.
MTS data typically have many channels with each variate corresponding to a channel. Many models, such as Informer [36], FEDformer [37], and Autoformer [29], handle MTS data to extract useful representations in a channel-mixing way, where the MTS input is treated as a two-dimensional matrix whose size is the number of channels multiplied by the length of history. Nonetheless, recently a few works such as PatchTST [23] and TiDE [6] have shown that a channel-independence way for handling MTS may achieve SOTA accuracy, where each channel is input to the model as a one-dimensional vector independent of the other channels. We believe that these two ways of handling LTSF need to be adopted as per the characteristics of the MTS data, rather than using a one-size-fits-all approach. When there are strong between-channel correlations, channel mixing usually can help capture such dependencies; otherwise, channel independence is a more sensible choice. Therefore, it is necessary to design a unified architecture applicable to both channel-mixing and channel-independence scenarios.
@@
In this paper, we introduce a novel approach that effectively captures long-range dependencies in time series data by providing sensible multi-scale contexts and particularly enhancing local contexts in the channel-independence situation. Our model, built upon a selective scan SSM called Mamba [11], serves as a core inference engine with a strong ability to capture long-range dependencies in MTS data while maintaining linear scalability and small memory footprints. The proposed model exploits the unique property of time series data in a bottom-up manner by producing contextual cues at two scales through consecutive resolution reduction or downsampling using linear mapping. The first level operates at a high resolution, while the second level works at a low resolution. At each level, we employ two Mamba modules to glean contextual cues from global perspectives for the channel-mixing case and from both global and local perspectives for the channel-independence case.
@@ 2 Related Works
Non-Transformer-based Supervised Approaches include classical methods like ARIMA, VARMAX, GARCH [5], and RNN [15], as well as deep learning-based methods that achieve state-of-the-art (SOTA) performance using multi-layer perceptrons (MLPs) and convolutional neural networks (CNNs). MLP-based models, such as DLinear [33], TiDE [6], and RLinear [18], leverage the simplicity of linear structures to achieve complexity and scalability.
CNN-based methods, such as TimesNet [30] and Scinet [19], utilize convolutional filters to extract valuable temporal features and model complex temporal dynamics. These approaches exhibit highly competitive performance, often comparable to or even occasionally outperforming more sophisticated Transformer-based models.
Transformer-based Supervised Learning Methods, such as iTransformer [21], PatchTST [23], Crossformer [35], FEDformer [37], stationary [20], and autoformer [29], have gained popularity for LTSF due to their superior accuracy. These methods convert time series to token sequences and leverage the self-attention mechanism to discover dependencies between arbitrary time steps, making them particularly effective for modeling complex temporal relationships. They may also exploit Transformers’ ability to process data in parallel, enabling long-term dependency discovery sometimes with even linear scalability. Despite their distinctive advantages, these methods typically have quadratic time and memory complexity due to point-wise correlations in self-attention mechanisms.
@@ 3 Proposed Method
Figure 1: Schematic diagram of our proposed methodology, TimeMachine. Our method incorporates a configuration of four Mambas, with two specialized Mambas capable of processing the transposed signal data in each branch. To the left, an example of the TimeSeries signals is depicted, while the right side offers a detailed zoomed-in view of a Mamba’s structure. Mambas are capable of accepting an input of shape $BMn_{i}$ while providing the same shape as the output. In our method $i\in\{1,2\}$.
@@
In this section, we describe each component of our proposed architecture and how we use our model to solve the LTSF problem. Assume a collection of MTS samples is given, denoted by dataset $\mathcal{D}$, which comprises an input sequence $x=[x_{1},\dots,x_{L}]$, with each $x_{t}\in{\mathcal{R}}^{M}$ representing a vector of $M$ measurements at time point $t$. The sequence length $L$ is also known as the look-back window. The goal is to predict $T$ future values, denoted by $[x_{L+1},\dots,x_{L+T}]$. The architecture of our proposed model, referred to as TimeMachine, is depicted in Figure 1. The pillars of this architecture consist of four Mambas, which are employed in an integrated way to tap contextual cues from MTS. This design choice enables us to harness Mamba’s robust capabilities of inferring sequential data for LTSF.
@@
Normalization: Before feeding the data to our model, we normalize the original MTS $x$ into $x^{0}=[x^{(0)}_{1},\cdots,x^{(0)}_{L}]\in{\mathcal{R}^{M\times L}}$, via $x^{(0)}=Normalize(x)$. Here, $Normalize(\cdot)$ represents a normalization operation with two different options. The first is to use the reversible instance normalization (RevIN) [16], which is also adopted in PatchTST [23]. The second option is to employ regular $Z$-score normalization: $x^{(0)}_{i,j}=({x_{i,j}-mean(x_{i,j}))/\sigma_{j}}$, where $\sigma_{j}$ is the standard deviation for channel $j$, with $j=1,\cdots,M$. Empirically we find that RevIN is often more helpful compared to Z-score. Apart from normalizing the data in the forward pass of our approach, in experiments, we also follow the standardization process of the data when compared with baseline methods.
@@
Channel Mixing vs. Channel Independence: Our model can handle both channel independence and channel mixing cases. In channel independence, each channel is processed independently by our model, while in channel mixing, the MTS sequence is processed with multiple channels combined throughout our architecture. Regardless of the case, our model accepts input of the shape $BML$ and produces the desired output of the shape $BMT$, eliminating the need for additional manual pre-processing.
@@
Channel independence has been proven effective in reducing overfitting by PatchTST [23]. We found this strategy helpful for datasets with a smaller number of channels. However, for datasets with a number of channels comparable to the look-back, channel mixing is more effective in capturing the correlations among channels and reaching the desired minimum loss during training.
@@
$x^{(1)}=E_{1}(x^{(0)}),\quad x^{(2)}=E_{2}(DO(x^{(1)})),$ (1)
@@
Integrated Quadruple Mambas: With the two processed embedded representations from $E_{1},E_{2}$, we can now learn more comprehensive representations by leveraging Mamba, a type of SSM with selective scan ability. At each embedding level, we employ a pair of Mambas to capture long-term dependencies within the look-back samples and provide sufficient local contexts. Denote the input to one of the four Mamba blocks by $u$, which is either $DO(x^{(1)})$ obtained after $E_{1}$ and the subsequent dropout layer for the two outer Mambas, or $DO(x^{(2)})$ obtained after $E_{2}$ and the subsequent dropout layer for the two inner Mambas (Figure 1). The input tensors may be reshaped per channel mixing or channel independence cases as described previously.
@@
where $h(t)$ is $N$-dimensional, with $N$ also known as a state expansion factor, $u(t)$ is $D$-dimensional, with $D$ being the dimension factor for an input token, $v(t)$ is an output of dimension $D$, and $A$, $B$, and $C$ are coefficient matrices of proper sizes. This dynamic system induces a discrete SSM governing state evolution and outputs given the input token sequence through time sampling at $\{k\Delta\}$ with a $\Delta$ time interval. This discrete SSM is
@@
Processed embedded representation with tensor size $BMn_{1}$ is transformed by outer Mambas, while that with $BMn_{2}$ is transformed by inner Mambas, as depicted in Figure 1. For the channel-mixing case, the whole univariate sequence of each channel is used as a token with dimension factor $n_{2}$ for the inner Mambas. The outputs from the left-side and right-side inner Mambas, $v_{L,k},v_{R,k}\in{\mathcal{R}^{n_{2}}}$, are element-wise added with $x^{(2)}_{k}$ to obtain $x^{(3)}_{k}$ for the $k$-th token, $k=1,\cdots,M$. That is, by denoting $v_{L}=[v_{L,1},\cdots,v_{L,M}]\in{\mathcal{R}^{M\times n_{2}}}$ and similarly $v_{R}\in{\mathcal{R}^{M\times n_{2}}}$, we have $x^{(3)}=v_{L}\bigoplus v_{R}\bigoplus x^{(2)}$, with $\bigoplus$ being element-wise addition. Then, $x^{(3)}$ is linearly mapped to $x^{(4)}$ with $P_{1}:x^{(3)}\rightarrow x^{(4)}\in{\mathcal{R}^{M\times n_{1}}}$. Similarly, the outputs from the outer Mambas, $v^{*}_{L,k},v^{*}_{R,k}\in{\mathcal{R}^{n_{1}}}$ are element-wise added to obtain $x^{(5)}\in{\mathcal{R}^{M\times n_{1}}}$.
@@
For the channel independence case, the input is reshaped, $BML\mapsto(B\times M)1L$, and the embedded representations become $(B\times M)1n_{1}$ and $(B\times M)1n_{2}$. One Mamba in each pair of outer Mambas or inner Mambas considers the input dimension as 1 and the token length as $n_{1}$ or $n_{2}$, while the other Mamba learns with input dimension $n_{2}$ or $n_{1}$ and token length 1. This design enables learning both global context and local context simultaneously. The outer and inner pairs of Mambas will extract salient features and context cues are fine and coarse scales with high- and low-resolution, respectively.
@@
Channel mixing is performed when the datasets contain a significantly large number of channels, in particular, when the look-back $L$ is comparable to the channel number $M$, taking the whole sequence as a token to better provide context cues. All four Mambas are used to capture the global context of the sequences at different scales and learn from the channel correlations. This helps stabilize the training and reduce overfitting with large $M$. To switch between the channel-independence and channel-mixing cases, the input sequence is simply transposed, with one Mamba in each branch processing the transposed input, as demonstrated in Figure 1.
These integrated Mamba blocks empower our model for content-dependent feature extraction and reasoning with long-range dependencies and feature interactions.
@@
Output Projection: After receiving the output tokens from the Mambas, our goal is to project these tokens to generate predictions with the desired sequence length. To accomplish this task, we utilize two MLPs, $P_{1}$ and $P_{2}$, which output $n_{1}$ and $T$ time points, respectively, with each point having $M$ channels. Specifically, projector $P_{1}$ performs a mapping $\mathcal{R}^{M\times n_{2}}\rightarrow\mathcal{R}^{M\times n_{1}}$, as discussed above for obtaining $x^{(4)}$. Subsequently, projector $P_{2}$ performs a mapping $\mathbb{R}^{M\times 2n_{1}}\rightarrow\mathbb{R}^{M\times T}$, transforming the concatenated output from the Mambas into the final predictions. The use of a two-stage output projection via $P_{1}$ and $P_{2}$ symmetrically aligns with the two-stage embedded representation obtained through $E_{1}$ and $E_{2}$.
@@ 4.1 Datasets
We evaluate our model on seven benchmark datasets extensively used for LTSF: Weather, Traffic, Electricity, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). Table 1 illustrates the relevant statistics of these datasets, highlighting that the Traffic and Electricity datasets notably large, with 862 and 321 channels, respectively, and tens of thousands of temporal points in each sequence. More details on these datasets can be found in Wu et al. [29], Zhou et al. [36]. Focusing on long-term forecasting, we exclude the ILI dataset, which has a shorter temporal horizon, similar to Das et al. [6]. We demonstrate the superiority of our model in two parts: quantitative (main results) and qualitative results. For a fair comparison, we used the code from PatchTST [23]  [^1] https://github.com/yuqinie98/PatchTST and iTransformer [21]  [^2] https://github.com/thuml/iTransformer, and we took the results for the baseline methods from iTransformer [21].
Table 1: Overview of the characteristics of used benchmarking datasets. Time points illustrate the total length of each dataset

### v2

@@ 1 Introduction
Capturing long-term dependencies, which are generally abundant in MTS data, is pivotal to LTSF. While linear models such as DLinear [34] and TiDE [6] achieve competitive performance with linear complexity and scalability, with accuracy on par with Transformer-based models, they usually rely on MLPs and linear projections that may not well capture long-range correlations [4]. Transformer-based models such as iTransformer [21], PatchTST [23], and Crossformer [36] have a strong ability to capture long-range dependencies and superior performance in LTSF accuracy, thanks to the self-attention mechanisms in Transformers [28]. However, they typically suffer from the quadratic complexity [6], limiting their scalability and applicability, e.g., in edge computing.
@@
The under-utilization of SSMs in LTSF can be attributed to two main reasons. First, highly content- and context-selective SSMs have only been recently developed [11]. Second, and more importantly, effectively representing the context in time series data remains a challenge. Many Transformer-based models, such as Autoformer [29] and Informer [37], regard each observation as a token in a sequence, while more recent models like PatchTST [23] and iTransformer [21] leverage patches of the time series as tokens. However, our empirical experiments on real-world MTS data suggest that directly utilizing SSMs for LTSF by using either observations or patches as tokens could hardly achieve performance comparable to Transformer-based models. Considering the particular characteristics of MTS data, it is essential to extract more salient contextual cues tailored to SSMs.
MTS data typically have many channels with each variate corresponding to a channel. Many models, such as Informer [37], FEDformer [38], and Autoformer [29], handle MTS data to extract useful representations in a channel-mixing way, where the MTS input is treated as a two-dimensional matrix whose size is the number of channels multiplied by the length of history. Nonetheless, recently a few works such as PatchTST [23] and TiDE [6] have shown that a channel-independence way for handling MTS may achieve state-of-the-art (SOTA) accuracy, where each channel is input to the model as a one-dimensional vector independent of the other channels. We believe that these two ways of handling LTSF need to be adopted as per the characteristics of the MTS data, rather than using a one-size-fits-all approach. When there are strong between-channel correlations, channel mixing usually can help capture such dependencies; otherwise, channel independence is a more sensible choice. Therefore, it is necessary to design a unified architecture applicable to both channel-mixing and channel-independence scenarios.
@@
In this paper, we introduce a novel approach that effectively captures long-range dependencies in time series data by providing rich multi-scale contexts and particularly enhancing local contexts in the channel-independence situation. Our model, built upon a selective scan SSM called Mamba [11], serves as a core inference engine with a strong ability to capture long-range dependencies in MTS data while maintaining linear scalability and small memory footprints. The proposed model exploits the unique property of time series data in a bottom-up manner by producing contextual cues at two scales through consecutive resolution reduction or downsampling using linear mapping. The first level operates at a high resolution, while the second level works at a low resolution. At each level, we employ two Mamba modules to glean contextual cues from global perspectives for the channel-mixing case and from both global and local perspectives for the channel-independence case.
@@ 2 Related Works
Non-Transformer-based Supervised Approaches include classical methods like ARIMA, VARMAX, GARCH [5], and RNN [15], as well as deep learning-based methods that achieve SOTA performance using multi-layer perceptrons (MLPs) and convolutional neural networks (CNNs). MLP-based models, such as DLinear [34], TiDE [6], and RLinear [18], leverage the simplicity of linear structures to achieve complexity and scalability. CNN-based methods, such as TimesNet [30] and Scinet [19], utilize convolutional filters to extract valuable temporal features and model complex temporal dynamics. These approaches exhibit highly competitive performance, often comparable to or even occasionally outperforming more sophisticated Transformer-based models.
Transformer-based Supervised Learning Methods, such as iTransformer [21], PatchTST [23], Crossformer [36], FEDformer [38], stationary [20], Flowformer [31], and Autoformer [29], have gained popularity for LTSF due to their superior accuracy. These methods convert time series to token sequences and leverage the self-attention mechanism to discover dependencies between arbitrary time steps, making them particularly effective for modeling complex temporal relationships. They may also exploit Transformers’ ability to process data in parallel, enabling long-term dependency discovery sometimes with even linear scalability. Despite their distinctive advantages, these methods typically have quadratic time and memory complexity due to point-wise correlations in self-attention mechanisms.
@@ 3 Proposed Method
Figure 1: Schematic diagram of our proposed model, TimeMachine. Our method incorporates a configuration of four Mambas, with two specialized Mambas capable of processing the transposed tensor data in each branch. On the left, an example of the MTS is depicted, while the right side shows a detailed view of a Mamba’s structure. Mambas are capable of accepting an input of shape $BMn_{i}$ while providing the output of the same shape, where $i\in\{1,2\}$ in our method.
@@
In this section, we describe each component of our proposed architecture and how we use our model to solve the LTSF problem. Assume a collection of MTS samples is given, denoted by dataset $\mathcal{D}$, which comprises an input sequence $\mathbf{x}=[x_{1},\dots,x_{L}]$, with each $x_{t}\in{\mathcal{R}}^{M}$ representing a vector of $M$ channels at time point $t$. For matrices, we use bold font; for scalars and vectors, we use regular (non-bold) letters. The sequence length $L$ is also known as the look-back window. The goal is to predict $T$ future values, denoted by $[x_{L+1},\dots,x_{L+T}]$. The architecture of our proposed model, referred to as TimeMachine, is depicted in Figure 1. The pillars of this architecture consist of four Mambas, which are employed in an integrated way to tap contextual cues from MTS. This design choice enables us to harness Mamba’s robust capabilities of inferring sequential data for LTSF.
@@
Normalization: Before feeding the data to our model, we normalize the original MTS $\mathbf{x}$ into $\mathbf{x}^{0}=[\mathbf{x}^{(0)}_{1},\cdots,\mathbf{x}^{(0)}_{L}]\in{\mathcal{R}^{M\times L}}$, via $\mathbf{x}^{(0)}=Normalize(\mathbf{x})$. Here, $Normalize(\cdot)$ represents a normalization operation with two different options. The first is to use the reversible instance normalization (RevIN) [16], which is also adopted in PatchTST [23]. The second option is to employ regular $Z$-score normalization: $x^{(0)}_{i,j}=({x_{i,j}-mean(x_{i,j}))/\sigma_{j}}$, where $\sigma_{j}$ is the standard deviation for channel $j$, with $j=1,\cdots,M$. Empirically we find that RevIN is often more helpful compared to Z-score. Apart from normalizing the data in the forward pass of our approach, in experiments, we also follow the standardization process of the data when compared with baseline methods.
@@
Channel Mixing vs. Channel Independence: Our model can handle both channel independence and channel mixing cases. In channel independence, each channel is processed independently by our model, while in channel mixing, the MTS sequence is processed with multiple channels combined throughout our architecture. Regardless of the case, our model accepts input of the shape $BML$, where $B$ is batch size, and produces the desired output of the shape $BMT$, eliminating the need for additional manual pre-processing.
@@
Channel independence has been proven effective in reducing overfitting by PatchTST [23]. We found this strategy helpful for datasets with a smaller number of channels. However, we observe for datasets with a number of channels comparable to the look-back, channel mixing is more effective in capturing the correlations among channels and reaching the desired minimum loss during training.
@@
$\mathbf{x}^{(1)}=E_{1}(\mathbf{x}^{(0)}),\quad\mathbf{x}^{(2)}=E_{2}(DO(\mathbf{x}^{(1)})),$ (1)
@@
Integrated Quadruple Mambas: With the two processed embedded representations from $E_{1},E_{2}$, we can now learn more comprehensive representations by leveraging Mamba, a type of SSM with selective scan ability. At each embedding level, we employ a pair of Mambas to capture long-term dependencies within the look-back samples and provide sufficient local contexts. Denote the input to one of the four Mamba blocks by $u$, which is either $DO(\mathbf{x}^{(1)})$ obtained after $E_{1}$ and the subsequent dropout layer for the two outer Mambas, or $DO(\mathbf{x}^{(2)})$ obtained after $E_{2}$ and the subsequent dropout layer for the two inner Mambas (Figure 1). The input tensors may be reshaped per channel mixing or channel independence cases as described.
@@
where $h(t)$ is $N$-dimensional, with $N$ also known as a state expansion factor, $u(t)$ is $D$-dimensional, with $D$ being the dimension factor for an input token, $v(t)$ is an output of dimension $D$, and $A$, $B$, and $C$ are coefficient matrices of proper sizes. This dynamic system induces a discrete SSM governing state evolution and outputs given the input token sequence through time sampling at $\{k\Delta\}$. Here, $\Delta$ is a time interval for discretizing the dynamic system. In particular, Mamba makes $\Delta$ a function of the input, and hence so do the model coefficients (A, B, C) and hidden state, thereby adapting the model dynamics to input and enhancing context selectivity. Consequently, this discrete SSM is
@@
Processed embedded representation with tensor size $BMn_{1}$ is transformed by outer Mambas, while that with $BMn_{2}$ is transformed by inner Mambas, as depicted in Figure 1. For the channel-mixing case, the whole univariate sequence of each channel is used as a token with dimension factor $n_{2}$ for the inner Mambas. The outputs from the left-side and right-side inner Mambas, $v_{L,k},v_{R,k}\in{\mathcal{R}^{n_{2}}}$, are element-wise added with $x^{(2)}_{k}$ to obtain $x^{(3)}_{k}$ for the $k$-th token, $k=1,\cdots,M$. That is, by denoting $\mathbf{v}_{L}=[v_{L,1},\cdots,v_{L,M}]^{T}\in{\mathcal{R}^{M\times n_{2}}}$ and similarly $\mathbf{v}_{R}\in{\mathcal{R}^{M\times n_{2}}}$, we have $\mathbf{x}^{(3)}=\mathbf{v}_{L}\bigoplus\mathbf{v}_{R}\bigoplus\mathbf{x}^{(2)}$, with $\bigoplus$ being element-wise addition. Then, $\mathbf{x}^{(3)}$ is linearly mapped to $\mathbf{x}^{(4)}$ with $P_{1}:\mathbf{x}^{(3)}\rightarrow\mathbf{x}^{(4)}\in{\mathcal{R}^{M\times n_{1}}}$. Similarly, the outputs from the outer Mambas, $v^{*}_{L,k},v^{*}_{R,k}\in{\mathcal{R}^{n_{1}}}$ are element-wise added to obtain $\mathbf{x}^{(5)}\in{\mathcal{R}^{M\times n_{1}}}$.
@@
For the channel independence case, the input is reshaped, $BML\mapsto(B\times M)1L$, and the embedded representations become $(B\times M)1n_{1}$ and $(B\times M)1n_{2}$. Here, the batch size becomes $B\times M$, and we regard each sequence of length $L$ independent from each other. One Mamba in each pair of outer Mambas or inner Mambas considers the input dimension as 1 and the token length as $n_{1}$ or $n_{2}$, while the other Mamba learns with input dimension $n_{2}$ or $n_{1}$ and token length 1. This design enables learning both global context and local context simultaneously. The outer and inner pairs of Mambas will extract salient features and context cues at fine and coarse scales with high- and low-resolution, respectively.
@@
Channel mixing is performed when the datasets contain a significantly large number of channels, in particular, when the look-back $L$ is comparable to the channel number $M$, taking the whole sequence as a token to better provide context cues. All four Mambas are used to capture the global context of the sequences at different scales and learn from the channel correlations. This helps stabilize the training and reduce overfitting with large $M$. To switch between the channel-independence and channel-mixing cases, the input sequence is simply transposed, with one Mamba in the outer Mamba pair and one in the inner pair processing the transposed input, as demonstrated in Figure 1. These integrated Mamba blocks empower our model for content-dependent feature extraction and reasoning with long-range dependencies and feature interactions.
@@
Output Projection: After receiving the output tokens from the Mambas, our goal is to project these tokens to generate predictions with the desired sequence length. To accomplish this task, we utilize two MLPs, $P_{1}$ and $P_{2}$, which output $n_{1}$ and $T$ time points, respectively, with each point having $M$ channels. Specifically, projector $P_{1}$ performs a mapping $\mathcal{R}^{M\times n_{2}}\rightarrow\mathcal{R}^{M\times n_{1}}$, as discussed above for obtaining $\mathbf{x}^{(4)}$. Subsequently, projector $P_{2}$ performs a mapping $\mathbb{R}^{M\times 2n_{1}}\rightarrow\mathbb{R}^{M\times T}$, transforming the concatenated output from the Mambas into the final predictions. The use of a two-stage output projection via $P_{1}$ and $P_{2}$ symmetrically aligns with the two-stage embedded representation obtained through $E_{1}$ and $E_{2}$.
@@ 4.1 Datasets
We evaluate our model on seven benchmark datasets extensively used for LTSF: Weather, Traffic, Electricity, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2). Table 1 illustrates the relevant statistics of these datasets, highlighting that the Traffic and Electricity datasets notably large, with 862 and 321 channels, respectively, and tens of thousands of temporal points in each sequence. More details on these datasets can be found in Wu et al. [29], Zhou et al. [37]. Focusing on long-term forecasting, we exclude the ILI dataset, which has a shorter temporal horizon, similar to Das et al. [6]. We demonstrate the superiority of our model in two parts: quantitative (main results) and qualitative results. For a fair comparison, we used the code from PatchTST [23]  [^1] https://github.com/yuqinie98/PatchTST and iTransformer [21]  [^2] https://github.com/thuml/iTransformer including normalization and evaluation protocol used by them, and we took the results for the baseline methods from iTransformer [21].
Table 1: Overview of the characteristics of used benchmarking datasets. Time points illustrate the total length of each dataset.